OpenStudy (anonymous):

Find the integral of x^2/(x^2-1) dx.

OpenStudy (dumbcow):

use polynomial division to simplify __1 R 1______ x^2 - 1 | x^2 -( x^2 -1) -------------- 1 \[\int\limits 1 + \frac{1}{x^2 - 1}\] use partial fractions

OpenStudy (anonymous):

I get it. Thanks for the help.

OpenStudy (dumbcow):

\[\frac{1}{(x-1)(x+1)} = \frac{A}{x-1} + \frac{B}{x+1}\] \[A + B = 0\] \[A -B = 1\] \[A = \frac{1}{2} , B = -\frac{1}{2}\] \[\int\limits dx +\frac{1}{2}\int\limits \frac{dx}{x-1} -\frac{1}{2}\int\limits \frac{dx}{x+1}\]